Database last updated on Sun, 01 Feb 2026 12:40:51 +0000 by localhost

Critical warnings from b-files' comments for days between 2024012 and 2024012 (0 sec, 5 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2024 012 18 07 19 Motor #1: appears stuck with a reference sensor activated.31B01224.191TORON#19105:28:271-1
rl:2024 012 18 08 17 New value set for watermark SUPPLY.VOLTAGE.HIGH[213B01224.191TORON#19105:28:341-1
skc:Azimuth drive turned OFF in op_st.12B01224.191TORON#19119:59:001-1
hg:Lamp not on... hg test terminated F1= 2 F4= 40921B01224.191TORON#19120:12:311-1
hg:Lamp not detected f1= 2 f4= 40921B01224.191TORON#19120:12:321-1

Instrument constant changes for the week before 2024012 (2 sec, 0 rows)

SL burned out? (AP results when SL is ON) (0 sec, 6 rows)

HG burned out? (AP results when HG is ON) (0 sec, 11 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
19120240121761017.0

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
19120240121139.4

Non-critical warnings from b-files' comments for days between 2024012 and 2024012 (1 sec, 7 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2024 012 18 07 19 Checksum bad in low level command.13B01224.191TORON#19105:28:251-1
rl:2024 012 18 07 19 Motor #1: appears stuck with a reference sensor activated.13B01224.191TORON#19105:28:271-1
rl:2024 012 18 07 19 Motor #3: appears stuck with a reference sensor activated.13B01224.191TORON#19105:28:301-1
rl:2024 012 18 07 19 Motor #11: appears stuck with a reference sensor activated.13B01224.191TORON#19105:28:321-1
hp:HOME key pressed499B01224.191TORON#19111:46:141-1
pd:Going to menu399B01224.191TORON#19111:46:461-1
hd:Illegal command (hd), number 9/ 14, in schedule , ignoring.199B01224.191TORON#19113:39:041-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssl
012TORONTO#191Azimuth drive turned OFF in op_st.,Going to menu,menu,Must do HG now,Saving backup OP_ST.000,Starting schedule inside,Waiting until ZA = 170 now 131.9,Waiting until ZA = 170 now 133.7,Waiting until ZA = 170 now 135.4,Waiting until ZA = 170 now 137.2,Waiting until ZA = 170 now 138.9,Waiting until ZA = 170 now 140.6,Waiting until ZA = 170 now 142.3,Waiting until ZA = 170 now 143.9,Waiting until ZA = 170 now 145.5,Waiting until ZA = 170 now 147.1,Waiting until ZA = 170 now 148.6,Waiting until ZA = 170 now 150,Waiting until ZA = 170 now 151.4,Waiting until ZA = 170 now 152.6,Waiting until ZA = 170 now 153.8,Waiting until ZA = 170 now 154.8,Waiting until ZA = 170 now 155.7,Waiting until ZA = 170 now 156.5,Waiting until ZA = 170 now 157.1,Waiting until ZA = 170 now 157.5,Waiting until ZA = 170 now 157.732282335
(1 sec)

Relative and absolute humidity for day 2024012 (1 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
19132.58.0